In triangle abc the vertices are A(-3,4), B(6,-2), C(2,5). Find the equation of the median passing through A.
Answer by ewatrrr(24785)


(-3,4)
(4, 3/2) Midpoint BC
m = (5/2)/-7 = - 5/14
y - 4 = (-5/14)(x + 3)
y = (-5/14)x + 41/14